OpenStudy (anonymous):

I tried to solve this and got 5 which isn't a option on the multiple choice... "Solve square root of the quantity two x plus three equals x plus one for x" A:2 B:+or-radical 2 C:radical 2 D:4 Please help!

OpenStudy (cwrw238):

sqrt(2x + 3) = x + 1 square both sides 2x + 3 = (x + 1)^2

OpenStudy (cwrw238):

2x + 3 = x^2 + 2x + 1 0 = x^2 - 2 can you finish this off?

OpenStudy (anonymous):

I don't think I can finish it. I don't understand these at all

OpenStudy (cwrw238):

x^2 = 2 now to get x you take the square root of both sides

OpenStudy (anonymous):

so the answer is A?

OpenStudy (cwrw238):

no x^2 = 2 square root of x^2 = x therefore x = +/- square root of 2 or as you say in the states +/- radical 2

OpenStudy (anonymous):

okay thank you!!!!

OpenStudy (cwrw238):

yw